OpenStudy (anonymous):

Simplify: 3(radical 12)−2(radical 75)+ (Radical 27)

OpenStudy (anonymous):

\[3\sqrt{12} -2\sqrt{75} + \sqrt{27} ?\]

OpenStudy (anonymous):

@Blairortiz

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[-\sqrt{3}\]

OpenStudy (anonymous):

How